• Use induction to prove for all n∈Nn∈N that n∑k=02k=2n+1−1.
1
Expert's answer
2021-06-16T17:26:25-0400

Prove "\displaystyle\sum_{k=0}^n2^k=2^{n+1}-1"


Let "P(n)" be the proposition that


"2^0+2^1+2^2+...2^n=2^{n+1}-1"

for the integer "n."


BASIS STEP:

"P(0)" is true because "2^0=1=2^{0+1}-1"

This completes the basis step.


INDUCTIVE STEP: For the inductive hypothesis, we assume that "P(k)" is true for an arbitrary nonnegative integer "k." That is, we assume that


"2^0+2^1+2^2+...2^k=2^{k+1}-1"

To carry out the inductive step using this assumption, we must show that when we assume that "P(k)" is true, then "P(k+1)" is also true. That is, we must show that


"2^0+2^1+2^2+...2^k+2^{k+1}=2^{k+1+1}-1=2^{k+2}-1"

assuming the inductive hypothesis "P(k)." Under the assumption of "P(k)," we see that


"2^0+2^1+2^2+...2^k+2^{k+1}"

"(2^0+2^1+2^2+...2^k)+2^{k+1}="

"=2^{k+1}-1+2^{k+1}"

"=2\cdot2^{k+1}-1"

"=2^{k+2}-1"

Note that we used the inductive hypothesis in the second equation in this string of equalities to replace "2^0+2^1+2^2+...2^k" by "2^{k+1}-1." We have completed the inductive step.

Because we have completed the basis step and the inductive step, by mathematical induction we know that "P(n)" is true for all nonnegative integers "n." That is,


"\displaystyle\sum_{k=0}^n2^k=2^{n+1}-1"

for all nonnegative integers "n."